# Normal forms for ordinary differential operators, I

### v3

@@ Abstract
In this paper we develop the generalised Schur theory offered in the recent paper by the second author in dimension one case, and apply it to obtain two applications in different directions of algebra/algebraic geometry.
The first application is a new explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups.
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators, which we prove in the case when operators have a normal form with the restriction top line (for details see Introduction).
Both applications are obtained with the help of normal forms. Namely, considering the ring of ordinary differential operators $D_{1}=K[[x]][\partial]$ as a subring of a certain complete non-commutative ring $\hat{D}_{1}^{sym}$, the normal forms of differential operators mentioned here are obtained after conjugation by some invertible operator (”Schur operator”), calculated using one of the operators in a ring. Normal forms of commuting operators are polynomials with constant coefficients in the differentiation, integration and shift operators, which have a restricted finite order in each variable, and can be effectively calculated for any given commuting operators.
@@ 1 Introduction
In this paper we develop the generalised Schur theory offered in [46] in dimension one case and apply it to obtain two applications in different directions of algebra/algebraic geometry.
The first application is a new convenient explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups. It is well known (see e.g. [5] or review [47]) that such sheaves are exactly the spectral sheaves of commutative subrings of ordinary differential operators of rank one. This result is the first step in establishing similar parametrisation for spectral sheaves of arbitrary rank, and also for spectral sheaves of commutative subrings of operators in higher dimensions (cf. [6], [45]). It is motivated by an important problem that appears in algebraic-geometric classification of commutative subrings of operators in higher dimensions – a description of torsion free sheaves with specific conditions on cohomology groups (see [45]), in particular with fixed Hilbert polynomial and some extra conditions. In the work [6] a description of Cohen-Macaulay sheaves on the spectral surface of quantum Calogero-Moser systems was given with the help of matrix problem approach due to I. Burban and Y. Drozd [3], [4] (Cohen-Macaulay sheaves form an open part of the moduli space of torsion free sheaves with fixed Hilbert polynomial, and by that reason it is important to describe them at first), however this approach meets with a difficulty to describe sheaves with specific cohomological properties (cf. [6, Sec. 6]). We expect that our new approach will help to solve this problem in an effective way in any dimension.
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators. It is motivated by the following natural question from the Burchnall-Chaundy theory.
The famous Burchnall-Chaundy lemma ([7]) says that any two commuting differential operators $P,Q\in D_{1}:=K[[x]][\partial]$ are algebraically dependent. More precisely, if the orders $n,m$ of operators $P,Q$ are coprime111i.e. the rank of the ring $K[P,Q]$ is 1, see e.g. [47] for relevant definitions, in particular [47, Lemma 5.23] for a proof of the Burchnall-Chaundy lemma in general form. The statement about the form of polynomial follows easily from the proof., then there exists an irreducible polynomial $f(X,Y)$ of weighted degree $v_{n,m}(f)=mn$ of special form (here the weighted degree is defined as in Dixmier’s paper [14], cf. item 4 in the List of Notations below): $f(X,Y)=\alpha X^{m}\pm Y^{n}+\ldots$ (here $\ldots$ mean terms of lower weighted degree, $0\neq\alpha\in K$; in particular, for coprime $p$ and $q$ the polynomial $f$ is automatically irreducible), such that $f(P,Q)=0$. A similar result for commuting operators of rank $r$ was established in [42] (cf. [35, Th. 2.11]), in this case $m=\mathop{\rm ord}\nolimits(Q)/r$, $n=\mathop{\rm ord}\nolimits(P)/r$, and again $GCD(m,n)=1$. Vice versa, if $P,Q\in D_{1}$ is a solution of such polynomial $f(X,Y)$ [^2] A solution of the equation $f(X,Y)=0$ is a pair $(P,Q)\in D$ such that $\sum_{i,j=0}^{n}\alpha_{ij}P^{i}Q^{j}=0$., then $[P,Q]=0$. Now a natural question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ for generic polynomial $F$ appears. This question appears to be surprisingly difficult in general case. We give a partial affirmative answer on this question in the case when the normal form has the restriction top line (see discussion below).
The necessity of further development of the Schur theory was not restricted only by the above mentioned applications. Let’s recall two major theorems of this theory from [46] (we formulate them here in a simplified form, for notations see the List of notations below):
@@
Assume $[\partial_{q}^{k},Q]=0$ for $q=1,\ldots,n$, where $Q\in\hat{D}_{n}^{sym}$. Then
@@
where all coefficients $c_{j_{1},\ldots,j_{n}}$ are polynomials in $\partial_{q}$, $\int_{q}$, $q=1,\ldots,n$ with constant coefficients and the degree of these polynomials with respect to $\partial_{q}$ is not greater than $\mathop{\rm\bf ord}(Q)$ and the degree of these polynomials with respect to $\int_{q}$ is not greater than $k-1$.
If the operators $P_{i}$ from the first theorem are differential, i.e. $P_{i}\in D_{n}$, a natural question appears: what is the shape of the operator $S$? In dimension one case it is well known (see e.g. [30], [22], [38] or the book [47] and references therein) that, if the centraliser of $P_{1}$ is non-trivial, then $S$ can be expressed via the Baker-Akhieser function (and vice versa), though in quite non-trivial way, if the rank of the centraliser is greater than one. Starting from works [18], [19], [20] it is known that the Baker-Akhieser function plays an important role in many problems of mathematical physics, in particular it played a key role in constructing explicit examples of commuting operators in many works (cf. [29], [28] for examples in dimension one, and [10], [11], [12] for examples in higher dimensions). Analogously, in higher dimension the Schur-Sato operator $S$ determines the common eigenfunction of commuting operators (of different nature), see [6, Sec. 6], [45], cf. also [2] and references therein, and the knowledge of its shape could help to prove the classification conjecture [45, Conj. 7.11] about commuting partial differential operators. Besides, in higher dimension, the operator $S$ determines an order-preserving endomorphism of the Weyl algebra ([46, Cor. 2.1]), thus giving hint to the Dixmier and Jacobian conjectures.
@@
To study normal forms of non-commuting operators we develop a technique of Newton regions (see section 4) – this is a natural generalisation of the technique of Newton polygons widely used for study of operators in the Weyl algebra (cf. [14], [17], [23]). Since normal forms of non-commuting operators are usually infinite series, the convex hull of all monomials may not be a restricted domain. However, in this case it is possible to define relevant notions of weights and top lines (generalisations of corresponding notions from [14]). In section 4 we study normal forms of a pair of non-commuting monic differential operators $P,Q\in D_{1}$. After conjugating this pair by a Schur operator of, say, operator $Q$, we obtain a monic operator $P^{\prime}\in\hat{D}_{1}^{sym}$ satisfying condition $A_{q}(0)$ (where $q$ is the order of $Q$). It is possible to define a weight function $v_{\sigma,\rho}$ and a notion of related top line for such operators. We distinguish 2 principal cases of top lines: the restriction top line and the asymptotic top line, both lines are uniquely defined (see definitions 4.3 and 4.4). Lemma 4.1 says that there are only two possibilities for a non commuting with $\partial^{q}$ operator $P^{\prime}$: it has either the restriction top line or the asymptotic top line. In section 4.3 we give the affirmative answer on the question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ in the case when the normal form $P^{\prime}$ of the pair $P,Q$ has the restriction top line.
We will consider the remaining case of the asymptotic top line as well as a similar description of the moduli space for sheaves of rank $>1$ in the next article, since these cases require much more detail. We expect that further study of normal forms (both for commuting and non-commuting operators) is reasonable not only for differential operators, but also for operators of other type, like difference, integro-differential, etc. In particular, it seems to be promising to study normal forms of recently discovered examples from [25], [24].
@@
In section 3 we study normal forms of commuting differential operators with the help of theory from section 2. We give a convenient description of the centraliser and of normal forms of a pair of commuting operators. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve.
In section 4 we study normal forms of non-commuting differential operators. In section 4.1 we introduce the notion of Newton region – a natural generalisation of the Newton polygon – for operators from $\hat{D}_{1}^{sym}$ and study its basic properties for operators satisfying condition $A_{k}(0)$ (all normal forms of differential operators satisfy this condition). In section 4.2 we prove one general combinatorial lemma, and in section 4.3 we prove the main theorem of section 4 – a commutativity criterion of a pair of operators in the case when the normal form of this pair has the restriction top line.
In Appendix we collect all necessary basic technical assertions about the weight function $v_{\sigma,\rho}$ and the homogeneous highest terms $f_{\sigma,\rho}$ used in section 4, with detailed proofs.
Acknowledgements. This work of the first author was partially supported by the National Key R and D Program of China (Grant No. 2020YFE0204200). The work of the second author was partially supported by RSF grant no. 22-11-00272.
@@ 1.1 List of notations
1. $K$ is a field of characteristic zero. Recall some notation from [46]: $\hat{R}:=K[[x_{1},\ldots,x_{n}]]$, the $K$-vector space
@@
$\sigma(P):=P_{\mathop{\rm\bf ord}(P)}=P_{-d}$ – the highest symbol.
2. In this paper we use: $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$,
@@
Operators: $\delta:=\exp((-x)\ast\partial)$, $\int:=(1-exp((-x)\ast\partial))\cdot\partial^{-1}$, $A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ (in the case when $k$ is fixed, simply written as $A_{i}$), where $\tilde{K}=K[\xi]$, $\xi$ is a primitive $k$th root of unity, $\Gamma_{i}=(x\partial)^{i}$. $B_{n}=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$.
@@
Operators written in the (Standard) form as
@@
are called HCP and form a sub-ring $Hcpc(k)$. They can be written also in G-form:
@@
$\Phi:\tilde{K}[A_{1},\ldots,A_{k-1}]\rightarrow\tilde{K}^{\oplus k},\quad P\mapsto(1\circ P,\ldots,\partial^{k-1}\circ P)$
@@
4. In section 4, suppose $H$ is an operator whose components are all HCP. Then $E(H)$ denotes the point set where $f_{l,i;r}\neq 0$, $v_{\sigma,\rho}$ stands for the weight degree of $H$, and $f_{\sigma,\rho}$ for the highest terms associated to $(\sigma,\rho)$:
$v_{\sigma,\rho}(H)=\sup\{\sigma l+\rho j|(l,j)\in E(H)\}\quad f_{\sigma,\rho}(H)=\sum_{(l,j)\in E(H,\sigma,\rho)}\sum_{i}f_{l,i;j}\Gamma_{l}A_{k,i}D^{j}$
The up-edge of the Newton region of $P$ is the set
$Edg_{u}(P):=\{(a,b)\in E(P)|\quad a=Sdeg_{A}(P_{b})\mbox{\quad and\quad}\forall b^{\prime}>b\quad Sdeg_{A}(P_{b^{\prime}})<a\}.$
And $H_{d;(\sigma,\rho)}(H),HS_{d;(\sigma,\rho)}^{m}(H)$ stands for
$H_{d;(\sigma,\rho)}(L):=\sum_{\sigma l+\rho j\geq d}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}A_{i}\partial^{j}$
and
$HS_{d;(\sigma,\rho)}^{m}(L):=\sum_{\sigma l+\rho j\geq d;l\leq m}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}\partial^{j}$
@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
Following the notations in [46], denote $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$, define the $K$-vector space
@@
${}.\mathop{\rm\bf ord}(P):=\sup\{k-v(a_{k})|\quad k\in\mathbb{N}_{0}\}\in\mathbb{Z}\cup\{\infty\}.$
@@
    Operators from $\hat{D}^{sym}_{1}$ can realise arbitrary endomorphisms of the $K$-algebra $\hat{R}$ which are continuous in the $\mathfrak{m}$-adic topology.
@@
The Dirac delta is given by the series $\delta:=exp((-x)\ast\partial):=1-x\partial+\frac{1}{2!}x^{2}\partial^{2}-\ldots$, which satisfies $\delta\circ f(x)=f(0)$, and the operator of integration is given by the series
@@
Note that $\int$ is only the right inverse of $\partial$, because $\partial\int=1$ and $\int\cdot\partial=((1-exp((-x)\ast\partial))\cdot\partial^{-1}\cdot\partial)=1-\delta$.
@@
    $\mathop{\rm\bf ord}(P\cdot Q)\leq\mathop{\rm\bf ord}(P)+\mathop{\rm\bf ord}(Q)$, and the equality holds if $\sigma(P)\cdot\sigma(Q)\neq 0$,
@@
    $\sigma(P\cdot Q)=\sigma(P)\cdot\sigma(Q)$, provided $\sigma(P)\cdot\sigma(Q)\neq 0$,
@@
In particular, the function $-\mathop{\rm\bf ord}$ determines a discrete pseudo-valuation on the ring $\hat{D}_{1}^{sym}$.
@@
The first part of lemma coincides with [46, Lem. 7.2]. The last assertion follows easily from the proof of [46, Lem. 7.2]. Namely, $A$ is a polynomial iff the infinite system of linear equations hold
@@
The following claim is a partial case of [46, Prop. 7.1].
@@
    $A_{i}A_{j}=A_{i+j}\quad\partial^{p}A_{i}=\xi^{pi}A_{i}\partial^{p}\quad A_{i}x^{p}=\xi^{pi}x^{p}A_{i}\quad\int^{p}A_{i}=\xi^{-pi}A_{i}\int^{p}.$
@@
    For a given $Q\in\hat{D}^{sym}_{1}$ assume that $[\partial^{k},Q]=0$.
@@
For convenience of the reader we’ll give here a proof of item 2 which is easier in our case that the proof of [46, Prop. 7.1] (and we’ll use some of its arguments later).
The identity $[\partial^{k},Q]=0$ can be rewritten as
@@
Any solution of the last equation in the commutative ring $\tilde{R}:=\tilde{K}[[x]]((\tilde{\partial}))$ has the form
$Q^{\prime}=c_{0}+c_{1}\exp((\xi_{k}-1)x\tilde{\partial})+\ldots+c_{k-1}\exp((\xi_{k}^{k-1}-1)x\tilde{\partial}),$
where $c_{i}\in\tilde{R}$ don’t depend on $x$ (as it follows from elementary differential algebra) [^3] First note that $\ker\{\partial:\tilde{R}\rightarrow\tilde{R}\}=\tilde{K}((\tilde{\partial}))$. Suppose there is a solution $H$ such that it is linearly independent with $1,\exp((\xi_{k}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$. Then $H_{0}:=\partial(H)=H^{\prime}$ is not equal to zero and is linearly independent with $\exp((\xi_{k}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$. By induction, $H_{i}:=(\partial+(1-\xi_{k}^{i})\tilde{\partial})(H_{i-1})=H_{i-1}^{\prime}+(1-\xi_{k}^{i})\tilde{\partial}H_{i-1}$ is not equal to zero and is linearly independent with $\exp((\xi_{k}^{i+1}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$ for all $i\leq k-1$, in particular $H_{k-1}\neq 0$. On the other hand, $H_{k-1}=\prod_{i=1}^{k}(\partial+(1-\xi_{k}^{i})\tilde{\partial})(H)=0$, a contradiction..
@@
For any $i=0,\ldots,k-1$ we have $\mathop{\rm\bf ord}(c_{i})\leq\mathop{\rm\bf ord}(Q)$ in formula (2.2), where the order $\mathop{\rm\bf ord}$ is defined on $\tilde{R}$ in the same way as on ^$\hat{{{\cal M}}}_{1}$.
@@
Since the elements $c_{i}$ are polynomials in ~$\tilde{\partial}^{-1}$, the expression in (2.2) can be written in the form
@@
where $\tilde{c}_{i}\in\tilde{K}[[\tilde{\partial}]]$ and $m\geq 0$, i.e. the series in brackets is divisible by ~$\tilde{\partial}^{m}$. We’ll additionally assume that $m$ is minimal, i.e. $GCD(\tilde{c}_{0},\ldots,\tilde{c}_{k-1})=1$ in the ring $\tilde{K}[[\tilde{\partial}]]$.
@@
Since $[\partial^{k},Q]=0$, we have also
@@
The expression $P_{(q)}$ is called a slice and the sum $P=\sum_{q\geq 0}P_{[q]}$ is called (a partial slice decomposition).
@@
(cf. [46], Def. 6.4) An element $P\in\hat{D}_{1}^{sym}$ is called regular if the $K$–linear map $F\xrightarrow{(-\circ\sigma(P))}F$ is injective, where $\circ$ means the action on the module $F$. In particular, $P$ is regular if and only if its symbol $\sigma(P)$ is regular.

### v4

@@ Abstract
In this paper we develop the generalised Schur theory offered in the recent paper by the second author in dimension one case, and apply it to obtain a new explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups.
This parametrisation is obtained with the help of normal forms - a notion we introduce in this paper. Namely, considering the ring of ordinary differential operators $D_{1}=K[[x]][\partial]$ as a subring of a certain complete non-commutative ring $\hat{D}_{1}^{sym}$, the normal forms of differential operators mentioned here are obtained after conjugation by some invertible operator (”Schur operator”), calculated using one of the operators in a ring. Normal forms of commuting operators are polynomials with constant coefficients in the differentiation, integration and shift operators, which have a restricted finite order in each variable, and can be effectively calculated for any given commuting operators.
@@ 1 Introduction
In this paper we develop the generalised Schur theory offered in [49] in dimension one case and apply it to obtain a new convenient explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups.
It is well known that such sheaves are exactly the spectral sheaves of commutative subrings of ordinary differential operators of rank one (see e.g. review [50] for detailed exposition of the theory of commuting ordinary differential operators or [5] for a short explanation). This result is the first step in establishing similar parametrisation for spectral sheaves of arbitrary rank, and also for spectral sheaves of commutative subrings of operators in higher dimensions (cf. [6], [48]). It is motivated by an important problem that appears in algebraic-geometric classification of commutative subrings of operators in higher dimensions – a description of torsion free sheaves with specific conditions on cohomology groups (see [48]), in particular with fixed Hilbert polynomial and some extra conditions. In the work [6] a description of Cohen-Macaulay sheaves on the spectral surface of quantum Calogero-Moser systems was given with the help of matrix problem approach due to I. Burban and Y. Drozd [3], [4] (Cohen-Macaulay sheaves form an open part of the moduli space of torsion free sheaves with fixed Hilbert polynomial, and by that reason it is important to describe them at first), however this approach meets with a difficulty to describe sheaves with specific cohomological properties (cf. [6, Sec. 6]). We expect that our new approach will help to solve this problem in an effective way in any dimension.
The necessity of further development of the Schur theory was not restricted only by the above mentioned application. Let’s recall two major theorems of this theory from [49] (we formulate them here in a simplified form, for notations see the List of notations below):
@@
Assume $\partial_{q}^{k},Q]=0$ for $q=1,\ldots,n$, where $Q\in\hat{D}_{n}^{sym}$. Then
@@
where all coefficients $c_{j_{1},\ldots,j_{n}}$ are polynomials in $\partial_{q}$, $\int_{q}$, $q=1,\ldots,n$ with constant coefficients and the degree of these polynomials with respect to $\partial_{q}$ is not greater than $\mathop{\rm\bf ord}(Q)$ and the degree of these polynomials with respect to $\int_{q}$ is not greater than k1k-1.
If the operators $P_{i}$ from the first theorem are differential, i.e. $P_{i}\in D_{n}$, a natural question appears: what is the shape of the operator $S$? In dimension one case it is well known (see e.g. [30], [23], [39] or the book [50] and references therein) that, if the centraliser of $P_{1}$ is non-trivial, then $S$ can be expressed via the Baker-Akhieser function (and vice versa), though in quite non-trivial way, if the rank of the centraliser is greater than one. Starting from works [19], [20], [21] it is known that the Baker-Akhieser function plays an important role in many problems of mathematical physics, in particular it played a key role in constructing explicit examples of commuting operators in many works (cf. [29], [28] for examples in dimension one, and [10], [11], [12] for examples in higher dimensions). Analogously, in higher dimension the Schur-Sato operator $S$ determines the common eigenfunction of commuting operators (of different nature), see [6, Sec. 6], [48], cf. also [2] and references therein, and the knowledge of its shape could help to prove the classification conjecture [48, Conj. 7.11] about commuting partial differential operators. Besides, in higher dimension, the operator $S$ determines an order-preserving endomorphism of the Weyl algebra ([49, Cor. 2.1]), thus giving hint to the Dixmier and Jacobian conjectures [^1] Occasionally, the technique of normal forms developed in this paper appeared to be very useful in proving the Dixmier conjecture for the first Weyl algebra offered in a recent preprint by the second author: A.B. Zheglov, The conjecture of Dixmier for the first Weyl algebra is true, https://arxiv.org/abs/2410.06959 . We hope that further development of the normal form technique together with effective methods of positive characteristic from papers [17], [18], [41], [42], used in proving the equivalence of the Dixmier conjectures and the Jacobian conjectures, will allow people to come closer to solving at least the $JC_{2}$ conjecture. .
@@
We will consider a similar description of the moduli space for sheaves of rank >1>1 in the next article, since this case requires much more details. We expect that further study of normal forms is reasonable not only for differential operators, but also for operators of other type, like difference, integro-differential, etc. In particular, it seems to be promising to study normal forms of recently discovered examples from [25], [24].
@@
In section 3 we study normal forms of commuting differential operators with the help of theory from section 2. We give a convenient description of the centraliser and of normal forms of a pair of commuting operators.
With the help of this description we prove our main result, theorem 3.3, where a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve is given. In particular, this result indicates that the moduli space of spectral sheaves of rank one is an affine open subscheme of the compactified Jacobian (cf. remark 3.8).
The paper (especially section 3) makes substantial use of the theory of commuting ODOs, which has a rich history, and its generalisations offered in a series of papers by the second author and other collaborators. To better understand the ideas behind this work, we recommend reading the background of the recent review [50] (especially sections 9-11) and articles [6], [49].
Acknowledgements. The work of J. Guo was partially supported by the National Key R and D Program of China (Grant No. 2020YFE0204200). The work of A. Zheglov was partially supported by RSF grant no. 25-11-00210.
@@
We are also grateful to the anonymous referees whose remarks allowed to improve the exposition.
@@ 1.1 List of notations
[…]
@@
$\sigma(P):=P_{\mathop{\rm\bf ord}(P)}$ – the highest symbol.
2. In this paper we use: $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}$$\partial$],
@@
Operators: $\delta:=\exp((-x)\ast\partial)$ [^2] Here and further $\ast$ in all exponentials means that we consider normalized Taylor power series, i.e. the powers of $x$ always stand on the left of powers of $\partial$, for example $\delta:=\exp((-x)\ast\partial)=\sum_{k=0}^{+\infty}\frac{(-1)^{k}}{k!}x^{k}\partial^{k}$., $\int:=(1-exp((-x)\ast\partial))\partial^{-1}$, $A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ (in the case when $k$ is fixed, simply written as $A_{i}$), where $\tilde{K}=K[\xi]$, $\xi$ is a primitive $k$th root of unity, $\Gamma_{i}=(x\partial)^{i}$. $B_{n}=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$.
$\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ means the same ring $\hat{D}^{sym}_{1}$, but defined over the base field $\tilde{K}$.
@@
We denote $D^{i}=\partial^{i}$ if $i\geq 0$ and $\int^{-i}$ if $i<0$. Operators written in the (Standard) form as
@@
are called HCP and form a sub-ring $Hcpc(k)$. Here $f_{i;r}(x,A_{k;i},\partial)$ is a polynomial of $x,A_{k;i},\partial$, $\mathop{\rm\bf ord}(f_{i;r})=0$, of the form
[…]
They can be written also in G-form:
@@
$\Phi:\tilde{K}[A_{1},\ldots,A_{k-1}]\rightarrow\tilde{K}^{\oplus k},\quad{P\mapsto(\sum_{i}p_{i}\xi^{i},\ldots,\sum_{i}p_{i}\xi^{i(k-1)})}$
@@
The classical ring of pseudo-differential operators is defined as
$E=K[[x]]((\partial^{-1})).$
There is an isomorphism of $\tilde{K}-$algebras $\psi:\mathfrak{B}\rightarrow M_{k}(C(\mathfrak{B}))$, where $C(\mathfrak{B})\simeq\tilde{K}[\tilde{D}^{k},\tilde{D}^{-k}]$, ($\tilde{K}$ is diagonally embedded into $\tilde{K}^{\oplus k}$):
with $\psi(D^{l})=T^{l}$, and extended by linearity. The map $\psi$ can be obviously extended to
$\psi:\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\hookrightarrow M_{k}(\tilde{K}((\tilde{D}^{-k}))).$
@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
Following the notations in [49], denote $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}$$\partial$], define the $K$-vector space
@@
$\mathop{\rm\bf ord}(P):=\sup\{k-v(a_{k})|\quad k\in\mathbb{N}_{0}\}\in\mathbb{Z}\cup\{\infty\}.$
@@
[…]
@@
The Dirac delta is given by the series $\delta:=exp((-x)\ast\partial):=1-x\partial+\frac{1}{2!}x^{2}\partial^{2}$-$\ldots$, which satisfies $\delta\circ f(x)=f(0)$ [^4] We denote by $\circ$ here and further the action of operators from $\hat{D}_{1}^{sym}$ on elements of $\hat{R}$, determined by the module structure from item 2 of 2.1. We denote by $\cdot$ the multiplication in the ring $\hat{D}_{1}^{sym}$. Sometimes the sign $\cdot$ is omitted, and in these cases we always mean $\cdot$, not $\circ$. , and the operator of integration is given by the series
@@
Note that $\int$ is only the right inverse of $\partial$, because $\partial\int=1$ and $\int\partial=((1-exp((-x)\ast\partial))\partial^{-1}\partial)=1-\delta$.
@@
    $\mathop{\rm\bf ord}(PQ)\leq\mathop{\rm\bf ord}(P)+\mathop{\rm\bf ord}(Q)$, and the equality holds if $\sigma(P)\sigma(Q)\neq 0$,
@@
    $\sigma(PQ)=\sigma(P)\sigma(Q)$, provided $\sigma(P)\sigma(Q)\neq 0$,
@@
In particular, the function $\mathop{\rm\bf ord}$ determines a discrete pseudo-valuation on the ring $\hat{D}_{1}^{sym}$.
@@
The first part of lemma coincides with [49, Lem. 7.2]. The last assertion follows easily from the proof of [49, Lem. 7.2]. Namely, $A$ is a polynomial iff the infinite system of linear equations hold
@@
The following claim is a partial case of [49, Prop. 7.1].
@@
    $A_{i}A_{j}=A_{i+j},\quad\partial^{p}A_{i}=\xi^{pi}A_{i}\partial^{p},\quad A_{i}x^{p}=\xi^{pi}x^{p}A_{i},\quad\int^{p}A_{i}=\xi^{-pi}A_{i}\int^{p}.$
@@
    For a given $Q\in\hat{D}^{sym}_{1}$ assume that $\partial^{k},Q]=0$.
@@
For convenience of the reader we’ll give here a proof of item 2 which is easier in our case than the proof of [49, Prop. 7.1] (and we’ll use some of its arguments later).
The identity $\partial^{k},Q]=0$ can be rewritten as
@@
Any solution of the last equation in the commutative ring $)\tilde{R}:=\tilde{K}[[x]]((\tilde{\partial}$)) has the form
[…]
@@
For any $i=0,\ldots,k-1$ we have $\mathop{\rm\bf ord}(c_{i})\leq\mathop{\rm\bf ord}(Q)$ in formula (2.2), where the order $\mathop{\rm\bf ord}$ is defined on $\tilde{R}$ in the same way as on ${{\cal M}}_{1}$.
@@
[…]
@@
where $\tilde{c}_{i}\in\tilde{K}$$\tilde{\partial}$]] and $m\geq 0$, i.e. the series in brackets is divisible by $\tilde{\partial}^{m}$. We’ll additionally assume that $m$ is minimal, i.e. $GCD(\tilde{c}_{0},\ldots,\tilde{c}_{k-1})=1$ in the ring $\tilde{K}$$\tilde{\partial}$]].
@@
Since $\partial^{k},Q]=0$, we have also
@@
The expression $P_{(q)}$ is called a slice and the sum $P=\sum_{q\geq 0}P_{[q]}$ is called a partial slice decomposition.
@@
(cf. [49], Def. 6.4) An element $P\in\hat{D}_{1}^{sym}$ is called regular if the KK–linear map $F\xrightarrow{(-\circ\sigma(P))}F$ is injective, where $\circ$ means the action on the module $F$. In particular, $P$ is regular if and only if its symbol $\sigma(P)$ is regular.